OpenStudy (anonymous):

What is the quotient: (3x2 + 4x – 15) ÷ (x + 3) ?

OpenStudy (amistre64):

\(\color{red}{\small\text{quotient}}\) is the number without its remainder right?

OpenStudy (anonymous):

3x-5

OpenStudy (amistre64):

\(\color{blue}{\small\text{it easier to do on paper for}}\) me \(\color{purple}{\small\text{:)}}\)

OpenStudy (anonymous):

wow how to write blue?, or red?

OpenStudy (amistre64):

\(\color{purple}{\small\text{just yell aot and it changes colors :)}}\)

OpenStudy (anonymous):

how?? yell aot?

OpenStudy (amistre64):

\(\color{green}{\small\text{sometimes you gotta be very quiet}}\); \(\color{red}{\small\text{but sometimes really loud}}\)

OpenStudy (anonymous):

sometimes you gotta be very quiet, but sometimes really loud

OpenStudy (anonymous):

no it does not..!!

OpenStudy (amistre64):

\(\color{orange}{\small\text{you}}\) \(\color{purple}{\small\text{didnt}}\) \(\color{green}{\small\text{say}}\) \(\color{red}{\small\text{it}}\) \(\color{brown}{\small\text{with}}\) \(\color{red}{\small\text{feeling}}\) :)

OpenStudy (anonymous):

lol ^^

OpenStudy (anonymous):

[color=red]i did[/color]

OpenStudy (anonymous):

I love MIT! here i said this with feeling :)

OpenStudy (anonymous):

omg just say me the code plz?

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

\(\color{purple}{\small\text{say something}}\)

OpenStudy (anonymous):

\(\color{pink}{\small\text{write something not say something amistre :) }}\)

OpenStudy (amistre64):

lol good :)

OpenStudy (amistre64):

\(\color{blue}{\small\text{its nice}}\) to colorize

OpenStudy (amistre64):

\(\color{brown}{\small\textbf{and bold font}}\)

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

\(\color{red}{\small\text{yes}}\)\(\color{orange}{\small\text{ it}}\)\(\color{pink}{\small\text{ is}}\)

OpenStudy (amistre64):

\(\color{red}{\small\text{only if you say please with feeling :)}}\)

OpenStudy (anonymous):

pleeeeeeeaaaaaaassssseeeeeee with feeling ;)

OpenStudy (anonymous):

well you wrote it, amistre wnated you to say not to write ^^

OpenStudy (amistre64):

\(\color{yellow}{\small\textbf{lol}}\)

OpenStudy (anonymous):

\(\color{pink}{\small\text{\[E(\phi , k) >\int\limits_{0}^{\phi}\sqrt{1-k^{2}\sin^{2}\theta} *d \theta\]}}\)

OpenStudy (anonymous):

\(\color{pink}{\small\text{ \[E(\phi , k) >\int\limits_{0}^{\phi}\sqrt{1-k^{2}\sin^{2}\theta} *d \theta\] }}\)

OpenStudy (anonymous):

well i can'T color equations :(

OpenStudy (amistre64):

\(\color{purple}{\frac{2}{3}}\)

OpenStudy (amistre64):

take out the \text part; thats only there for words

OpenStudy (amistre64):

\[\color{purple}{\sqrt{18}}\]

OpenStudy (anonymous):

(\color{pink}{\small{ \[E(\phi , k) >\int\limits_{0}^{\phi}\sqrt{1-k^{2}\sin^{2}\theta} *d \theta\] }}\)

OpenStudy (amistre64):

\[\color{orange}{2\pi r^2}\]

OpenStudy (amistre64):

\[ \color{green} {E (\phi , k) > \int \limits_{0}^{ \phi } \sqrt{1-k^{2}\sin^{2}\theta} *d \theta\}}\]